# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ A.1. Omitted definitions
${}\wedge\forall\bar{x}\Big{(}\phi_{\mathfrak{A}}(\bar{x})\vee\bigvee\nolimits_{x_{1},x_{2}\in\bar{x}}\neg E(x_{1},x_{2})\Big{)}.$
@@
A.6. A full proof of Theorem 2.2 and Theorem 2.4
@@
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(\mathbb{Q};<)\wr\mathfrak{B}$. By Example 1.3 and Example 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. We then define $(\mathbb{Q};<)^{\mathfrak{B}}$ as the expansion of $\mathfrak{B}^{\prime}$ by the binary inequality $\neq$. Note that $(\mathbb{Q};<)^{\mathfrak{B}}$ has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $(\mathbb{Q};<)^{\mathfrak{B}}_{I_{4}}$. We obtain the structure $\mathbb{Q}^{\mathfrak{A}}$ as a reduct of $(\mathbb{Q};<)^{\mathfrak{B}}$ by removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$; but note that $\mathbb{Q}^{\mathfrak{A}}$ can also be defined directly from $\mathfrak{A}$. We continue with the proofs of CSP/Pol-injectivity and Datalog-interreducibility, starting with the CSP-injective case.
@@
We claim that $\mathbb{Q}^{\mathfrak{A}}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow\mathbb{Q}^{\mathfrak{A}}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{\mathbb{Q}^{\mathfrak{A}}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$. Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $\mathbb{Q}^{\mathfrak{A}}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $\mathbb{Q}^{\mathfrak{A}}$ is CSP-injective.
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation.
@@
It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $\mathbb{Q}^{\mathfrak{A}}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$: otherwise, the map on $\mathfrak{A}$ sending every element to $h(q_{\sim}(x))$ would provide such constant endomorphism. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $\mathbb{Q}^{\mathfrak{A}}$. By the CSP-injectivity of $\mathbb{Q}^{\mathfrak{A}}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $\mathbb{Q}^{\mathfrak{A}}$.
We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow\mathbb{Q}^{\mathfrak{A}}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{\mathbb{Q}^{\mathfrak{A}}}$ restricted to $i(C)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow\mathbb{Q}^{\mathfrak{A}}$.
@@
We now consider $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$. By Proposition 4.3, $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ is Pol-injective and $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}}_{I_{4}})$ and $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}})$ are Datalog-interreducible. That $\mathrm{CSP}(\mathbb{Q}^{\mathfrak{A}}_{I_{4}})$ and $\mathrm{CSP}(\mathfrak{A}$) are Datalog-interreducible then follows from the fact that Datalog-reductions can be composed.
@@
Suppose that $e^{\ast}$ is an endomorphism of $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$, and let $S^{\ast}$ be an arbitrary finite subset of ×$\mathbb{Q}\times A$. Since $e^{\ast}$ preserves $\neq$, it is injective. Since $e^{\ast}$ preserves $E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}$, which is an equivalence relation, it naturally acts on its classes. Let $e$ be the map on (×$(\mathbb{Q}\times A)/E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}$ mapping each $[x]_{\smash{E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}}}$ to $[e^{\ast}(x)]_{\smash{E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}}}$, and let $S\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{[x]_{E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}}\mid x\in S^{\ast}\}$. Similarly as in Proposition 4.5(1), we have that $e$ is an endomorphism of $\mathfrak{A}$. Since $\mathfrak{A}$ is a model-complete core, there exists an automorphism $\alpha$ of $\mathfrak{A}$ such that $\alpha|_{S}=e|_{S}$. We claim that there exists an automorphism ${\alpha}^{\ast}$ of $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. To see this, note that we can simply fix a permutation of each equivalence class of $E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}$ and extend $\alpha$, which is a permutation of the equivalence classes, according to it. Since $S^{\ast}$ is finite, we can moreover select the permutations so that ${\alpha}^{\ast}$ coincides with $e^{\ast}$ on $S^{\ast}$.
@@
Next, suppose that $e$ is an endomorphism of $\mathfrak{A}$, and let $S$ be an arbitrary finite subset of $A$. Define $S^{\ast}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\{0\}\times S$. Let $e^{\ast}$ be the map on ×$\mathbb{Q}\times A$ defined via $e^{\ast}(q,a)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(q,e(a))$. This function is an endomorphism of $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$. Since $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ is a model-complete core, there exists an automorphism ${\alpha}^{\ast}$ of $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. Let $\alpha$ be the map on $A$ mapping each $[x]_{E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}}$ to $[e^{\ast}(x)]_{E^{\smash{\mathbb{Q}^{\mathfrak{A}}_{I_{4}}}}}$ As before, we have that $\alpha$ is a well-defined automorphism of $\mathfrak{A}$ extending $e$ on $S$.
@@
Proof of item 2. Since $(\mathbb{Q};<)^{\mathfrak{B}}_{I_{4}}$ is an expansion of $\mathfrak{B}^{\prime}$ by two relations that are preserved by all embeddings, we can ignore these relations and only prove the statement for $\mathfrak{B}^{\prime}$. Recall that $\mathfrak{B}^{\prime}=(\mathbb{Q};<)\wr\mathfrak{B}$. The first part of item 2 follows directly from Proposition 4.5(4) because $(\mathbb{Q};<)$ is homogeneous. We continue with the second part. There the backward direction follows directly from Proposition 4.5(6) because $(\mathbb{Q};<)$ is homogeneous Ramsey [54]. We continue with the forward direction, which we prove by hand.
@@
Suppose that $\mathrm{age}((\mathbb{Q};<)^{\mathfrak{B}}_{I_{4}})$ has the RP. Then clearly also $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP. Let $\mathfrak{X},\mathfrak{Y}\in\mathrm{age}(\mathfrak{B})$ be arbitrary. We transform them into $\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ by letting $E$ interpret as the diagonal relation and $<$ as the empty relation. Since $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP, there exists a Ramsey witness $\mathfrak{W}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$. Set $\mathfrak{W}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(\mathfrak{W}^{\prime}/E^{\mathfrak{W}^{\prime}})|^{\sigma}$. Let $f\colon\smash{\binom{\mathfrak{W}}{\mathfrak{X}}}\rightarrow[2]$ be arbitrary. We define $f^{\prime}\colon\smash{\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}}\rightarrow[2]$ as follows. For an embedding $e^{\prime}\in\smash{\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}}$, we set $f^{\prime}(e^{\prime})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=f(q_{E}\circ e^{\prime})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$. We have that $q_{E}\circ e^{\prime}\in\smash{\binom{\mathfrak{W}}{\mathfrak{X}}}$ because $E$ interprets as the diagonal relation in $\mathfrak{X}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$. Since $\mathfrak{W}^{\prime}$ is a Ramsey witness for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$, there exists $e^{\prime}\in\smash{\binom{\mathfrak{W}^{\prime}}{\mathfrak{Y}^{\prime}}}$ such that $f^{\prime}$ is constant on $\bigl{\{}e^{\prime}\circ u^{\prime}\mid u^{\prime}\in\smash{\binom{\mathfrak{Y}^{\prime}}{\mathfrak{X}^{\prime}}}\bigr{\}}\subseteq\smash{\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}}$. Since $E$ interprets as the diagonal relation in $\mathfrak{Y}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$, the map $e\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=q_{E}\circ e^{\prime}$ is contained in $\smash{\binom{\mathfrak{W}}{\mathfrak{Y}}}$. By construction, $f$ is constant on $\bigl{\{}e\circ u\;|\;u\in\smash{\binom{\mathfrak{Y}}{\mathfrak{X}}}\bigr{\}}\subseteq\smash{\binom{\mathfrak{W}}{\mathfrak{X}}}$.
@@
Proof of item 3. Homogeneity follows directly from the previous item. The finite boundedness can be deduced from Proposition 4.5(5), and the fact that we take an expansion by relations definable by a Boolean combination of equality atoms. More specifically, as explained in the introduction, up to isomorphism, the finite substructures of a finitely bounded structures are just the finite models of a universal sentence. By adding new clauses, that define the relations $\neq$ and $I_{4}$ in terms of equalities, to the universal sentence for $\mathrm{age}(\mathfrak{B}^{\prime})$, we obtain a universal sentence for $\mathrm{age}((\mathbb{Q};<)^{\mathfrak{B}})$.
@@
Proof of item 4. One direction is simple (and does not use the assumption on the orbit growth), because $\mathbb{Q}^{\mathfrak{A}}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=\smash{(\mathbb{Q}^{\mathfrak{A}}/E^{\mathbb{Q}^{\mathfrak{A}}})|^{\tau}}$. Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $\mathbb{Q}^{\mathfrak{A}}$.

### v2

@@ A.1. Omitted definitions
${}\wedge\forall\bar{x}\Big{(}\phi_{\mathfrak{A}}(\bar{x})\vee\bigvee\nolimits_{x_{1},x_{2}\in\bar{x}}\neg E(x_{1},x_{2})\Big{)}\Big{)}.$
@@
A.7. A full proof of Theorem 2.2 and Theorem 2.4
@@
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. By Example 1.3 and Example 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$ from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the proofs of CSP/Pol-injectivity and Datalog-interreducibility, starting with the CSP-injective case.
@@
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in ×$\mathbb{Q}\times\{a\}$ since this does not affect the relations in $\tau\cup\{E\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
@@
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$: otherwise, the map on $\mathfrak{A}$ sending every element to $h(q_{\sim}(x))$ would provide such constant endomorphism. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
We now consider $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. By Proposition 4.3, $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is Pol-injective and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ are Datalog-interreducible. That $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}})$ and $\mathrm{CSP}(\mathfrak{A}$) are Datalog-interreducible then follows from the fact that Datalog-reductions can be composed.
@@
Suppose that $e^{\ast}$ is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Let $S^{\ast}$ be an arbitrary finite subset of ×$\mathbb{Q}\times A$. Since $e^{\ast}$ preserves $\neq$, it is injective. Since $e^{\ast}$ preserves $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$, which is an equivalence relation, it naturally acts on its classes. Let $e$ be the map on (×$(\mathbb{Q}\times A)/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ mapping each $[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $[e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$, and let $S\coloneqq\{[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}\mid x\in S^{\ast}\}$. Similarly as in Proposition 4.5(1), we have that $e$ is an endomorphism of $\mathfrak{A}$. Since $\mathfrak{A}$ is a model-complete core, there exists an automorphism $\alpha$ of $\mathfrak{A}$ such that $\alpha|_{S}=e|_{S}$. We claim that there exists an automorphism ${\alpha}^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. To see this, note that we can simply fix a permutation of each equivalence class of $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}$ and extend $\alpha$, which is a permutation of the equivalence classes, according to it. Since $S^{\ast}$ is finite, we can moreover select the permutations so that ${\alpha}^{\ast}$ coincides with $e^{\ast}$ on $S^{\ast}$.
@@
Next, suppose that $e$ is an endomorphism of $\mathfrak{A}$, and let $S$ be an arbitrary finite subset of $A$. Define $S^{\ast}\coloneqq\{0\}\times S$. Let $e^{\ast}$ be the map on ×$\mathbb{Q}\times A$ defined via $e^{\ast}(q,a)\coloneqq(q,e(a))$. This function is an endomorphism of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ is a model-complete core, there exists an automorphism ${\alpha}^{\ast}$ of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ such that ${\alpha}^{\ast}|_{S^{\ast}}=e^{\ast}|_{S^{\ast}}$. Let $\alpha$ be the map on $A$ mapping each $[x]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ to $[e^{\ast}(x)]_{E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}}}$ As before, we have that $\alpha$ is a well-defined automorphism of $\mathfrak{A}$ extending $e$ on $S$.
@@
Proof of item 2. Since $\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}}$ is an expansion of $\mathfrak{B}^{\prime}$ by two relations that are preserved by all embeddings, we can ignore these relations and only prove the statement for $\mathfrak{B}^{\prime}$. Recall that $\mathfrak{B}^{\prime}=(\mathbb{Q};<)\wr\mathfrak{B}$. The first part of item 2 follows directly from Proposition 4.5(4) because $(\mathbb{Q};<)$ is homogeneous. We continue with the second part. There the backward direction follows directly from Proposition 4.5(6) because $(\mathbb{Q};<)$ is homogeneous Ramsey [52]. We continue with the forward direction, which we prove by hand.
@@
Suppose that $\mathrm{age}(\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq,I_{4}})$ has the RP. Then clearly also $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP. Let $\mathfrak{X},\mathfrak{Y}\in\mathrm{age}(\mathfrak{B})$ be arbitrary. We transform them into $\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ by letting $E$ interpret as the diagonal relation and $<$ as the empty relation. Since $\mathrm{age}(\mathfrak{B}^{\prime})$ has the RP, there exists a Ramsey witness $\mathfrak{W}^{\prime}\in\mathrm{age}(\mathfrak{B}^{\prime})$ for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$. We set $\mathfrak{W}\coloneqq(\mathfrak{W}^{\prime}/E^{\mathfrak{W}^{\prime}})|^{\sigma}$. Let $f\colon\binom{\mathfrak{W}}{\mathfrak{X}}\rightarrow[2]$ be arbitrary. We define $f^{\prime}\colon\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}\rightarrow[2]$ as follows. For an embedding $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$, we set $f^{\prime}(e^{\prime})\coloneqq f(q_{E}\circ e^{\prime})$, where $q_{E}$ denotes the factor map $x\mapsto[x]_{E}$. We have that $q_{E}\circ e^{\prime}\in\binom{\mathfrak{W}}{\mathfrak{X}}$ because $E$ interprets as the diagonal relation in $\mathfrak{X}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$. Since $\mathfrak{W}^{\prime}$ is a Ramsey witness for $(\mathfrak{X}^{\prime},\mathfrak{Y}^{\prime})$, there exists $e^{\prime}\in\binom{\mathfrak{W}^{\prime}}{\mathfrak{Y}^{\prime}}$ such that $f^{\prime}$ is constant on $\bigl{\{}e^{\prime}\circ u^{\prime}\;|\;u^{\prime}\in\binom{\mathfrak{Y}^{\prime}}{\mathfrak{X}^{\prime}}\bigr{\}}\subseteq\binom{\mathfrak{W}^{\prime}}{\mathfrak{X}^{\prime}}$. Since $E$ interprets as the diagonal relation in $\mathfrak{Y}^{\prime}$ and as a relational congruence w.r.t. $\sigma$ in $\mathfrak{W}^{\prime}$, the map $e\coloneqq q_{E}\circ e^{\prime}$ is contained in $\binom{\mathfrak{W}}{\mathfrak{Y}}$. By construction, $f$ is constant on $\bigl{\{}e\circ u\;|\;u\in\binom{\mathfrak{Y}}{\mathfrak{X}}\bigr{\}}\subseteq\binom{\mathfrak{W}}{\mathfrak{X}}$.
@@
Proof of item 3. The finite boundedness can be deduced from Proposition 4.5(5), and the fact that we take an expansion by relations definable by a Boolean combination of equality atoms. More specifically, as explained in the introduction, up to isomorphism, the finite substructures of a finitely bounded structures are just the finite models of a universal sentence. By adding new clauses, that define the relations $\neq$ and $I_{4}$ in terms of equalities, to the universal sentence for $\mathrm{age}(\mathfrak{B}^{\prime})$, we obtain a universal sentence for $\mathrm{age}(\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq})$.
@@
Proof of item 4. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=((\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}})|^{\tau}$. Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.